A consulting company is retained by a large organization for a 4-year project. The consulting fees in the 1st year will be $430,000. Costs are expected to decrease by 9% per year due to the learning curve and the consulting company's ability to work more quickly as it better understands the organization's procedures. The consulting company will bill the organization for the entire cost of the project at the end of year 4 and will charge 12% interest per year on consulting fees incurred. How much will the organization owe the consulting company at the end of year 4?
Answer #1

Solution :

1st Year Cost $430,000 plus Interest @ 12% is 51600(430000*12%) for one year & for 4 years it will be $ 206400(51600*4) since interest charged per year on consulting fees as stated.

So, Amount charged for 1st Year will be Cost+interest: $430,000+$206,400= $636400

2nd Year Cost will decrease by 9% as compared to the 1st Year, which can be calculated as follows:

1st Year Cost @ 430,000 * 91%= $391,300

2nd Year Interest will be calculated as follows:

Cost: 391300 * 12%= 46,956 ,

46956 * 3 = $140,868

So, Amount charged for 2nd Year will be Cost+interest: $391,300+$140,868 = $532,168

3rd Year Cost will decrease by 9% as compared to the 2nd Year, which can be calculated as follows:

2nd Year Cost @ 391300 * 91%= $356,083

3rd Year Interest will be calculated as follows:

Cost: 356083 * 12%= 42,730 ,

42730 * 2 = $85,460

So, Amount charged for 3rd Year will be Cost+interest: $356083+$85460 = $441543

4th Year Cost will decrease by 9% as compared to the 3rd Year, which can be calculated as follows:

3rd Year Cost @ 356083 * 91%= $324036

4th Year Interest will be calculated as follows:

Cost: 324036 * 12%= $38,884

So, Amount charged for 4th Year will be Cost+interest: $324036+$38884 = $362920

by adding all the amounts chared for all years we will get the answer:

$1,973,031 (1st Year:$636400 + 2nd year:$532,168 + 3rd year:$441543 + 4th year:$362920)

so, the organization owe the consulting company at the end of year 4 will be $1,973,031